# Adaptive operator learning for infinite-dimensional Bayesian inverse problems

### v2

@@ Abstract
The fundamental computational issues in Bayesian inverse problems (BIP) governed by partial differential equations (PDEs) stem from the requirement of repeated forward model evaluations. A popular strategy to reduce such costs is to replace expensive model simulations with computationally efficient approximations using operator learning, motivated by recent progress in deep learning. However, using the approximated model directly may introduce a modeling error, exacerbating the already ill-posedness of inverse problems. Thus, balancing between accuracy and efficiency is essential for the effective implementation of such approaches. To this end, we develop an adaptive operator learning framework that can reduce modeling error gradually by forcing the surrogate to be accurate in local areas. This is accomplished by adaptively fine-tuning the pre-trained approximate model with training points chosen by a greedy algorithm during the posterior computational process. To validate our approach, we use DeepOnet to construct the surrogate and unscented Kalman inversion (UKI) to approximate the BIP solution, respectively. Furthermore, we present a rigorous convergence guarantee in the linear case using the UKI framework. The approach is tested on a number of benchmarks, including the Darcy flow, the heat source inversion problem, and the reaction-diffusion problem. The numerical results show that our method can significantly reduce computational costs while maintaining inversion accuracy.
@@ 1 Introduction
Many realistic phenomenons are governed by partial differential equations (PDEs), where the states of the system are described by PDEs solutions. The properties of these systems are characterized by the model parameters, such as the permeability and thermal conductivity, which can not be directly determined. Instead, the parameters can be inferred from the discrete and noisy observations of the states, which are known as inverse problems. Because inverse problems are ill-posed in general, many methods for solving them are based primarily on either regularization theory or Bayesian inference. By imposing a prior distribution on the parameters, the Bayesian approach can provide a more flexible framework. The solutions to the Bayesian inverse problems, or the posterior distributions, can then be obtained by conditioning the observations using Baye’s formula. In some cases, the model parameters are required to be functions, leading to infinite-dimensional Bayesian inverse problems. These cases possibly occur when the model parameters are spatially varying with uncertain spatial structures, which can be found in many realistic applications, including many areas of engineering, sciences and medicine [1, 2, 3, 4, 5].
@@
The formulation of infinite-dimensional Bayesian inverse problems presents a number of challenges, including the well-posedness guaranteed by the proper prior selection, as well as the convergence of the solutions governed by the discretization scheme. Following that, dealing with the discrete finite-dimensional posterior distributions can be difficult due to the expensive-to-solve forward models and high-dimensional parameter spaces. As a result, direct sampling methods such as MCMC-based methods [6, 7, 8] will suffer from unaffordable computation costs. Common methods to deal with these issues include (i) model reduction methods [9, 1, 10, 11], which exploit the intrinsic low dimensionality, (ii) direct posterior approximation methods, such as Laplace approximation and variational inference [12, 4], and (iii) surrogate modeling [13, 14, 15], which replaces the expensive model with a cheap substitute.
@@
Surrogate modeling emerges as the most promising approach for efficiently accelerating the sampling of posterior distributions among the methods listed above. Deep learning methods, specifically deep neural networks (DNN), have recently become the most popular surrogate models in engineering and science due to their power in approximating high-dimensional problems[16, 17, 18, 19, 20]. In general, DNN employs the power of machine learning to construct a quick-to-evaluate surrogate model to approximate the parameter-to-observation maps[21, 15, 22]. Numerical experiments, such as those described in [15], demonstrated that with sufficiently large training datasets, highly accurate approximations can be trained. Traditional deep learning methods, on the other hand, frequently necessitate a lot of training points that are not always available. Furthermore, whenever the measurement operator changes, the surrogate should be retrained. Physical-informed neural networks(PINNs)[17] can address this issue by incorporating the physical laws into the loss function and learning the parameter-to-state maps[23, 24]. Due to this, they can be applied as surrogates for a class of various Bayesian inverse problems with models that are governed by the same PDEs but have various types of observations and noise modes, further reducing the cost of surrogate construction. However, PINNs have some limitations[25, 26], such as hyperparameter sensitivity and the potential for training instability due to the hybrid nature of their loss function. Several solutions have been proposed to address these issues[27, 28, 29, 30]. Furthermore, due to the requirement of a large collocation dataset, PINNs will continue to be ineffective for infinite-dimensional Bayesian inverse problems[23]. Operator neural networks, such as FNO[31] and DeepOnet[32], are able to model complex systems in high-dimensional spaces as infinite-dimensional approximations. They are therefore promising surrogates, as described in[33, 34]. However, using approximate models directly may introduce a discrepancy or modeling error, exacerbating an already ill-posed problem and leading to a worse solution.
We propose in this paper to develop a framework that can adaptively reduce model error by forcing the approximate model to be locally accurate for posterior characterization during the inversion process. This is achieved by first using neural network representations of parameter-to-state maps between function spaces, and then retraining this initial model with points chosen adaptively using a greedy algorithm. The inversion process can continue using the fine-tuned approximate model with lower local model error. This procedure can be repeated multiple times as necessary until the stop criteria is met. For the detailed implementation, we use the DeepOnet[32] to approximate the parameter-to-state maps and the unscented Kalman inversion [35] to estimate the posterior distribution. Moreover, we can show that under the linear case, convergence can be obtained if the surrogate is accurate throughout the space, which can also be extended to non-linear cases with locally accurate approximate models. To demonstrate the effectiveness of our method, we propose testing several benchmarks such as the Darcy flow, the heat source inversion problem, and a reaction-diffusion problem. Our main contributions can be summarized as follows.
@@ 2.1 Infinite-dimensional Bayesian inverse problems
$\begin{cases}&\mathcal{A}(u(\mathbf{x});m(\mathbf{x}))=0,\quad\mathbf{x}\in\Omega,\\ &\mathcal{B}(u(\mathbf{x}))=0,\quad\mathbf{x}\in\partial\Omega,\end{cases}$ (1)
where $\mathcal{A}$ denotes the general partial differential operator defined in the domain $\Omega\in\mathbb{R}^{d}$, $\mathcal{B}$ is the boundary operator on the boundary $\partial\Omega$, $m\in\mathcal{M}$ represents the unknown parameter field and $u\in\mathcal{U}$ represents the state field of the system.
Let $y\in\mathbb{R}^{N_{y}}$ denote a set of discrete and noisy observations at specific locations in $\Omega$. Suppose the state $u$ and $y$ are connected through an observation system $\mathcal{O}:\mathcal{U}\rightarrow\mathbb{R}^{N_{y}}$,
@@
$\nu(dm)=\frac{1}{Z(y)}\exp(-\Phi(m;y))\nu_{0}(dm),$ (4)
@@
$Z(y):=\int_{\mathcal{M}}\exp(-\Phi(m;y))\nu_{0}(dm).$
@@ 2.2 DeepOnet as surrogates
In this section, we employ the neural operator DeepOnet as the surrogate, which is fast to evaluate and can speed up in the posterior computational. The basic idea is to approximate the true forward operator $\mathcal{F}$ with a neural network $\mathcal{F}_{\theta}:\mathcal{M}\rightarrow\mathcal{U}$, where $\mathcal{M},\mathcal{U}$ are spaces defined before and $\theta$ are the parameters in the neural network. This neural operator can be interpreted as a combination of encoder $\mathcal{E}$, approximator $\mathcal{A}$ and reconstructor $\mathcal{R}$ [39] as depicted in Figs.1 and 2, i.e.,
@@
$\theta^{*}=\arg\min_{\theta\in\Theta}\mathcal{L}(\theta)=\int_{\mathcal{M}}\int_{\Omega}|\mathcal{F}(m)(\mathbf{x})-\mathcal{F}_{\theta}(m)(\mathbf{x})|^{2}\,\mathrm{d}\mathbf{x}\,\nu_{0}(dm),$ (5)
@@
Fig. 1: Diagram of the three components of DeepOnet.
Fig. 2: The framework of DeepOnet.
@@
$\widehat{\nu}(dm)\propto\exp(-\widehat{\Phi}(m;y))\nu_{0}(dm),$ (7)
@@
$\nu(\mathcal{M}(\epsilon))=\int_{\mathcal{M}(\epsilon)}\nu(dm).$
@@ 3.1 Adaptive model error reduction
Fig. 3: The flowchart of our framework.
While using sufficient training points to train the surrogate model $\widehat{\mathcal{G}}$ is not practical and will lose the advantage of computational efficiency [15, 40] especially for high dimensional problems. Instead, according to Theorem 2.1, the approximate posterior will be close to the true posterior when the surrogate is accurate in the high density region of the posterior distribution. This implies that we only need to maintain the local accuracy since the posterior distribution is typically concentrated in high dimensional spaces. However, the high density region of the posterior distribution is unknown until the observations are given. Therefore, we can design an algorithm that can update the training dataset during the posterior computational process to fine-tune the pre-trained surrogate. This refined surrogate can maintain the local accuracy as well as saving computational cost. Based on this purpose, we separate the adaptive algorithms into offline and online stages. Offline, a small amount of samples from the prior distribution are utilized to train the DeepOnet in an acceptable length of time. The purpose of offline computation is to obtain a rough but comprehensive approximation of the DeepOnet model. The online stage consists of two major steps. First, decide whether to adaptively update the surrogate model with the posterior computation algorithm. Second, if the surrogate model needs to be updated, choose a new training set locally. Our new approach is an online strategy that depends on a specific realization of the observed data and an associated posterior approximation, in contrast to the direct DeepOnet strategy which maintains the surrogate model $\widehat{\mathcal{G}}$ unchanged. We expect that the inversion results computed by adaptive DeepOnet will have better accuracy than those produced by the offline direct DeepOnet strategies because using the data centers attention on regions of high posterior probability and causes a localization effect in the construction of surrogate. We will explore this conjecture in numerical results below.
@@
The rest of this part turns this outline into a workable algorithm by explaining when and where to refine, as well as how to choose new training data to improve approximations. To create an initialization surrogate, we can start by creating a small training dataset $\mathcal{D}$ from the prior distribution. In detail, suppose we have a collection of model evaluations $\mathcal{D}=\{(m_{i},\mathcal{F}(m_{i}))\}$. We then can use these points to train an operator network $\mathcal{F}_{\theta}$ and obtain a surrogate $\widehat{\mathcal{G}}=\mathcal{O}\circ\mathcal{F}_{\theta}$, which can be evaluated repeatedly at negligible cost. Subsequently, we decide where and when to refine the surrogate by conducting exploration using the current surrogate. Specifically, assume that $\widehat{\nu}_{t}$ is the approximate posterior at $t$ step. We can now obtain $T$ samples $\mathbf{M}^{(t)}:=\{m_{k}^{(t)}\}_{k=1}^{T}$ from $\widehat{\nu}_{t}$ using a posterior computation method, such as particle methods or MCMC-based sampling methods. Next, we must choose a criteria to determine whether the surrogate need to be refined. Notice that our primary goal is to maintain the surrogate accurate along the posterior computation trajectories. A natural idea is to use this local model error as an indicator: if the local model error exceeds a predefined tolerance, the surrogate model need to refine. In other words, we have to reduce the local model error $e_{\mathcal{M}}$, which is defined by
$e_{\mathcal{M}}(t):=\mathbb{E}\left[\mathcal{G}-\mathcal{\widehat{G}}_{t}\right]=\left(\int_{\mathcal{M}}|\mathcal{G}(m)-\mathcal{\widehat{G}}_{t}(m)|^{2}\widehat{\nu}_{t}(dm)\right)^{\frac{1}{2}},$ (8)
as much as possible during the process of posterior computation. However, it will be prohibitively expensive to evaluate the local model error (8) in high-dimensional spaces. To determine whether we need to refine the surrogate, we use the least-squares functional with true forward defined in Eq.(3) as an indicator in this work. This is due to the fact that this value may indicate the discrepancy between the observed and predicted data, which may also indicate the accuracy of the inversion. To this end, we first define an anchor point $r^{t}$ from the obtained samples $\mathbf{M}^{(t)}$ using the following optimization:
@@
Then, the error indicator $e_{\mathcal{D}}(t)$ can be obtained as
@@
With a tolerance of $\epsilon$, if $|\frac{e_{\mathcal{D}}(t)-e_{\mathcal{D}}(t-1)}{e_{\mathcal{D}}(t)}|<\epsilon$, then the refinement process stops. If not, the surrogate must to refine in $\widehat{\nu}_{t}$.
@@
The following task is to develop sampling strategies for surrogate refinement. We only need to ensure that the surrogate is accurate along the posterior computational trajectories, so we can generate a small set of important (adaptive) points near the anchor points and then refine the surrogate. This ensures that the refined surrogate $\widehat{\mathcal{G}}_{t}$ is locally accurate in $\widehat{\nu}_{t}$ while also reducing computational costs. To this end, we propose to combine the numerical techniques used in posterior computation and a greedy algorithm to adaptively choose the most important samples from $\widehat{\nu}_{t}$. In order to cover the parameter space, we first draw a large set of $K$ samples $\Gamma=\{m_{1},m_{2},\cdots,m_{K}\}$ from $\widehat{\nu}_{t}$. A Greedy algorithm is used to select a subset $\gamma^{Q}=\{\widetilde{m}_{1},\widetilde{m}_{2},\cdots,\widetilde{m}_{Q}\}\subset\Gamma$ of “important” points one by one from $\Gamma$. That is, the newly selected point $\widetilde{m}_{j}$ must be close to the anchor point $r^{t}$. In addition, the surrogate solution for this point should have the greatest distance with the space spanned by $\{\widehat{\mathcal{G}}_{t}(\widetilde{m}_{i})\}_{i=1}^{j-1}$, i.e.,
$\widetilde{m}_{j}=\arg\max_{m\in\Gamma/\gamma^{j-1}}\left(D\big{(}\widehat{\mathcal{G}}_{t}(m),\widehat{\mathcal{G}}_{t}(\gamma^{j-1})\big{)}-d(m,r_{t})\right),$ (11)

### v3

@@ Abstract
The fundamental computational issues in Bayesian inverse problems (BIP) governed by partial differential equations (PDEs) stem from the requirement of repeated forward model evaluations. A popular strategy to reduce such costs is to replace expensive model simulations with computationally efficient approximations using operator learning, motivated by recent progress in deep learning. However, using the approximated model directly may introduce a modeling error, exacerbating the already ill-posedness of inverse problems. Thus, balancing between accuracy and efficiency is essential for the effective implementation of such approaches. To this end, we develop an adaptive operator learning framework that can reduce modeling error gradually by forcing the surrogate to be accurate in local areas. This is accomplished by adaptively fine-tuning the pre-trained approximate model with training points chosen by a greedy algorithm during the posterior evaluation process. To validate our approach, we use DeepOnet to construct the surrogate and unscented Kalman inversion (UKI) to approximate the BIP solution, respectively. Furthermore, we present a rigorous convergence guarantee in the linear case using the UKI framework. The approach is tested on a number of benchmarks, including the Darcy flow, the heat source inversion problem, and the reaction-diffusion problem. The numerical results show that our method can significantly reduce computational costs while maintaining inversion accuracy.
@@ 1 Introduction
Many realistic phenomenons are governed by partial differential equations (PDEs), where the states of the system are described by PDEs solutions. The properties of these systems are characterized by the model parameters, such as the permeability and thermal conductivity, which can not be directly determined. Instead, the parameters can be inferred from the discrete and noisy observations of the states, which are known as inverse problems. Because inverse problems are ill-posed in general, many methods for solving them are based primarily on either regularization theory or Bayesian inference. By imposing a prior distribution on the parameters, the Bayesian approach can provide a more flexible framework. The solutions to the Bayesian inverse problems, or the posterior distributions, can then be obtained by conditioning the observations using Baye’s formula. In some cases, the model parameters are required to be functions, leading to infinite-dimensional Bayesian inverse problems. These cases possibly occur when the model parameters are spatially varying with uncertain spatial structures, which can be found in many realistic applications, including many areas of engineering and science [1, 2, 3, 4, 5].
@@
The formulation of infinite-dimensional Bayesian inverse problems presents a number of challenges, including the well-posedness guaranteed by the proper prior selection, as well as the convergence of the solutions governed by the discretization scheme. Following that, dealing with the discrete finite-dimensional posterior distributions can be difficult due to the expensive-to-solve forward models and high-dimensional parameter spaces. Common methods to deal with these issues include (i) model reduction methods [6, 1, 7], which exploit the intrinsic low dimensionality of the governing physical systems, (ii) direct posterior approximation methods, such as Laplace approximation and variational inference [8, 4], and (iii) surrogate modeling [9, 10, 11, 12, 13], which approximates the computationally expensive model with a more efficient, lower-cost alternative.
@@
Surrogate modeling emerges as the most promising approach for efficiently accelerating the sampling of posterior distributions among the methods listed above. Deep learning methods, specifically deep neural networks (DNN), have recently become the most popular surrogate models in engineering and science due to their power in approximating high-dimensional problems[14, 15, 16, 17, 18]. In general, DNN employs the power of machine learning to construct a quick-to-evaluate surrogate model to approximate the parameter-to-observation maps[19, 13, 20]. Numerical experiments, such as those described in [13], demonstrated that with sufficiently large training datasets, highly accurate approximations can be trained. Traditional deep learning methods, on the other hand, frequently necessitate a lot of training points that are not always available. Furthermore, whenever the measurement operator changes, the surrogate should be retrained. Physical-informed neural networks(PINNs)[15] can address this issue by incorporating the physical laws into the loss function and learning the parameter-to-state maps[21, 22]. As a result, they can be used as surrogates for a variety of Bayesian inverse problems involving models governed by the same PDEs but with different types of observations operators, which will further reduce the cost of surrogate construction. However, PINNs have some limitations[23, 24], such as hyperparameter sensitivity and the potential for training instability due to the hybrid nature of their loss function. Several solutions have been proposed to address these issues[25, 26, 27, 28, 29]. Operator neural networks, such as FNO[30] and DeepOnet[31], are able to model complex systems in high-dimensional spaces as infinite-dimensional approximations. They are therefore promising surrogates, as described in[32, 33]. However, using approximate models directly may introduce a discrepancy or modeling error, exacerbating an already ill-posed problem and leading to a worse solution.
In order to reduce model errors, several approaches[9, 10, 13, 20, 34, 35] that incorporate local approximation techniques have been applied to Bayesian posterior sampling problems. For example, Conrad et al.[9] described a framework for building and refining local approximation during an MCMC simulation using either local polynomial approximations or local Gaussian process regressors(GPR). To generate the sample sets used for local approximations, the authors employ a sequential experimental design procedure that interleaves infinite refinement of the approximation with Markov chain posterior exploration. Recently, Cleary et al. [34] proposed a “Calibrate-Emulate-Sample”(CES) framework for approximate Bayesian inversion. This approach consists of three main steps: To begin, the ensemble Kalman method[36, 37] is used in conjunction with a full-order model to extract sample points from the exact posterior distribution. Next, these sample points are used to create a GPR emulator for the parameter-to-observation map. Finally, the approximated posterior obtained from the emulator is sampled using direct sampling methods such as MCMC. A similar idea is proposed in [20], which use a goal-oriented DNN surrogate approach to significantly reduce the computational burden of posterior sampling. Specifically, they begin by using a Laplace approximation to approximate the posterior distribution. After that, they select training points to build a DNN surrogate from this approximate distribution. In the last stage, direct sampling methods are used to sample from the approximate posterior. However, these works primarily focus on building local approximations of the parameter-to-observation maps. One significant limitation of this approach is that the calibrate phase (i.e., generating the posterior approximation) necessitates extensive computations with full-order models, which can be computationally expensive. Furthermore, if the observation data changes, the entire process must be restarted from scratch, including recalculating the posterior approximation and reconstructing the local approximation. This can reduce the framework’s efficiency and increase its resource requirements, especially in dynamic environments where observations change frequently. Inspired by these works, we present a mutual learning framework that can reduce model error by forcing the approximate model to be locally accurate for posterior characterization during the inversion process. This is achieved by first using neural network representations of parameter-to-state maps between function spaces, and then fine-tuning this initial model with points chosen adaptively from the approximate posterior. This procedure can be repeated multiple times as necessary until the stop criteria is met. In contrast to the CES framework, we choose training points from the prior distribution to build the initial emulator offline. When measurement data is available, the surrogate model only needs to be fine-tuned during the sampling process, which involves specific sampling methods. The integration of prior knowledge into the initial model construction enhances the robustness and accuracy of the model, while the targeted fine-tuning during sampling improves computational efficiency, making the approach particularly suitable for scenarios requiring fast and reliable decision-making. For the detailed implementation, we use the DeepOnet[31] to approximate the parameter-to-state maps and the unscented Kalman inversion [38] to estimate the posterior distribution. Moreover, we can show that under the linear case, convergence can be obtained if the surrogate is accurate throughout the space, which can also be extended to non-linear cases with locally accurate approximate models. To demonstrate the effectiveness of our method, we propose testing several benchmarks such as the Darcy flow, the heat source inversion problem, and a reaction-diffusion problem. Our main contributions can be summarized as follows.
@@ 2.1 Infinite-dimensional Bayesian inverse problems
$\begin{cases}&\mathcal{H}(u(\mathbf{x});m(\mathbf{x}))=0,\quad\mathbf{x}\in\Omega,\\ &\mathcal{B}(u(\mathbf{x});m(\mathbf{x}))=0,\quad\mathbf{x}\in\partial\Omega,\end{cases}$ (1)
where $\mathcal{H}$ denotes the general partial differential operator defined in the domain $\Omega\in\mathbb{R}^{d}$, $\mathcal{B}$ is the boundary operator on the boundary $\partial\Omega$, $m\in\mathcal{M}$ represents the unknown parameter field and $u\in\mathcal{U}$ represents the state field of the system.
Let $y\in\mathbb{R}^{N_{y}}$ denotes a set of discrete and noisy observations at specific locations in $\Omega$. Suppose the state $u$ and $y$ are connected through an observation system $\mathcal{O}:\mathcal{U}\rightarrow\mathbb{R}^{N_{y}}$,
@@
$\nu({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}m})=\frac{1}{Z(y)}\exp(-\Phi(m;y))\nu_{0}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}m}),$ (4)
@@
$Z(y):=\int_{\mathcal{M}}\exp(-\Phi(m;y))\nu_{0}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}m}).$
@@ 2.2 DeepOnet as surrogates
In this section, we employ the neural operator DeepOnet as the surrogate, which is fast to evaluate and can speed up in the posterior evaluations. The basic idea is to approximate the true forward operator $\mathcal{F}$ with a neural network $\mathcal{F}_{\theta}:\mathcal{M}\rightarrow\mathcal{U}$, where $\mathcal{M},\mathcal{U}$ are spaces defined before and $\theta$ are the parameters in the neural network. This neural operator can be interpreted as a combination of encoder $\mathcal{E}$, approximator $\mathcal{A}$ and reconstructor $\mathcal{R}$ [42] as depicted in Figs.1 and 2, i.e.,
@@
$\theta^{*}=\arg\min_{\theta\in\Theta}\mathcal{L}(\theta)=\int_{\mathcal{M}}\int_{\Omega}|\mathcal{F}(m)(\mathbf{x})-\mathcal{F}_{\theta}(m)(\mathbf{x})|^{2}\,\mathrm{d}\mathbf{x}\,\nu_{0}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}m}),$ (5)
@@
Figure 1: Diagram of the three components of DeepOnet.
Figure 2: The framework of DeepOnet.
@@
$\widehat{\nu}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}m})\propto\exp(-\widehat{\Phi}(m;y))\nu_{0}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}m}),$ (7)
@@
$\nu(\mathcal{M}(\epsilon))=\int_{\mathcal{M}(\epsilon)}\nu({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}m}).$
@@ 3.1 Adaptive model error reduction
Developing stable and reliable adaptive surrogate modeling methods presents several challenges, especially when dealing with infinite-dimensional Bayesian inverse problems. One major challenge lies in the need to maintain the accuracy of the surrogate model in the high-density regions of the posterior distribution, where the true posterior is most concentrated. According to Theorem 2.1, the approximate posterior will be close to the true posterior when the surrogate is accurate in the high density region of the posterior distribution. On the other hand, the accuracy of the surrogate model, especially in the context of an operator network like DeepOnet, depends heavily on the quality and distribution of the training set. Ideally, this training set should be well-distributed across the posterior distribution in order to capture the key characteristics of the problem. However, the high density region of the posterior distribution is unknown until the observations are given. One feasible approach, as discussed earlier, is the CES framework [34], which first involves a calibration step where posterior samples are obtained using the full-order model, and then these samples can be used to train the surrogate. Nonetheless, this strategy can be very computationally expensive, especially when new observational data emerges, as the entire process must be repeated. A natural question is how to create an effective local approximation that maintains a good balance between accuracy and computational cost. To achieve this, the new adaptive framework should be designed to automatically select new training points during the posterior exploration process. These selected points are then used to fine-tune the emulator generated by DeepOnet. Instead of relying on the full-order model to explore the posterior space initially, the training points for the emulator are sequentially updated by integrating samplers that utilize the emulator itself. Based on this purpose, we separate the adaptive algorithms into offline and online stages. Offline, a small amount of samples from the prior distribution are utilized to train the DeepOnet in an acceptable length of time. The purpose of offline computation is to obtain a rough but comprehensive approximation of the DeepOnet model. The online stage consists of two major steps. First, decide whether to adaptively update the surrogate model with the posterior computation algorithm. Second, if the surrogate model needs to be updated, choose a new training set locally. Our new approach is an online strategy that depends on a specific realization of the observed data and an associated posterior approximation, in contrast to the direct DeepOnet strategy which maintains the surrogate model $\widehat{\mathcal{G}}$ unchanged. We expect that the inversion results computed by adaptive DeepOnet will have better accuracy than those produced by the offline direct DeepOnet strategies because using the data centers attention on regions of high posterior probability and causes a localization effect in the construction of surrogate. We will explore this conjecture in numerical results below.
Figure 3: The flowchart for our mutual learning framework. The framework begins with an offline, pre-trained operator network. The operator network combines with observations to efficiently return likelihood information for samplers to explore the parameter space, while the samplers provide useful candidate points for surrogate refinement. They can learn from and improve each other over time.
@@
The rest of this part turns this outline into a workable algorithm by explaining when and where to refine, as well as how to choose new training data to improve approximations. To create an initialization surrogate, we can start by creating a small training dataset $\mathcal{D}$ from the prior distribution. In detail, suppose we have a collection of model evaluations $\mathcal{D}=\{(m_{i},\mathcal{F}(m_{i}))\}$. We then can use these points to train an operator network $\mathcal{F}_{\theta}$ and obtain a surrogate $\widehat{\mathcal{G}}=\mathcal{O}\circ\mathcal{F}_{\theta}$, which can be evaluated repeatedly at negligible cost. Subsequently, we decide where and when to refine the surrogate by conducting exploration using the current surrogate. Specifically, assume that $\widehat{\nu}_{t}$ is the approximate posterior at $t$ step. To assess the accuracy of the surrogate model in the approximate posterior space, we define the following “local” model error assist with $\widehat{\nu}_{t}$:
$e_{\mathcal{M}}(t):=\mathbb{E}\left[\mathcal{G}-\mathcal{\widehat{G}}_{t}\right]=\left(\int_{\mathcal{M}}|\mathcal{G}(m)-\mathcal{\widehat{G}}_{t}(m)|^{2}\widehat{\nu}_{t}({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mathrm{d}m})\right)^{\frac{1}{2}}.$ (8)
A natural approach is to use this error as an indicator: if the error exceeds a predefined tolerance, the surrogate model should be refined. Unfortunately, the need for high-dimensional integration makes directly evaluating this error in high-dimensional spaces prohibitively expensive. Notice that our primary goal is to maintain the surrogate accurate along the posterior computation trajectories. To achieve this, we can now obtain $T$ samples $\mathbf{M}^{(t)}:=\{m_{k}^{(t)}\}_{k=1}^{T}$ from $\widehat{\nu}_{t}$ using a posterior computation method, such as particle methods or MCMC-based sampling methods. Importantly, this process only relies on information from the surrogate model $\mathcal{\widehat{G}}_{t}$ and the data $y$, without requiring any information from the full-order model $\mathcal{G}$. To prevent the current samples from deviating too far from the true posterior trajectory, we define an anchor point $r^{t}$ from the obtained samples $\mathbf{M}^{(t)}$ using the full-order model $\mathcal{G}$ as follows:
@@
This allows us to define an error indicator $e_{\mathcal{D}}(t)$ based on the data-fitting term at the anchor point $r^{t}$ within the posterior sample set:
@@
This error indicator not only measures the model error but also ensures that the samples do not deviate significantly from the true posterior trajectory. When the relative error $|\frac{e_{\mathcal{D}}(t)-e_{\mathcal{D}}(t-1)}{e_{\mathcal{D}}(t)}|$ exceeds a tolerance of $\epsilon$, the surrogate must be refined in $\widehat{\nu}_{t}$. Otherwise, the refinement process stops.
@@
The following task is to develop sampling strategies for surrogate refinement. We only need to ensure that the surrogate is accurate along the posterior evaluation trajectories, so we can generate a small set of important (adaptive) points near the anchor point and then refine the surrogate. This ensures that the refined surrogate $\widehat{\mathcal{G}}_{t}$ is locally accurate in $\widehat{\nu}_{t}$ while also reducing computational costs. This topic is relevant to optimal experimental design. A number of existing strategies can be used to accomplish this goal. In this work, we propose a greedy algorithm for efficiently selecting the most important samples from $\widehat{\nu}_{t}$. Specifically, we first draw a large set of $K$ samples $\Gamma=\{m_{1},m_{2},\cdots,m_{K}\}$ from $\widehat{\nu}_{t}$. A greedy algorithm is then used to select a subset $\gamma^{Q}=\{\widetilde{m}_{1},\widetilde{m}_{2},\cdots,\widetilde{m}_{Q}\}\subset\Gamma$ of “important” points one by one from $\Gamma$. Assuming the current selected point sets are $\gamma^{j}(j<Q)$, then the newly selected point $\widetilde{m}_{j+1}$ must be near the anchor point $r^{t}$. Furthermore, the surrogate solution for this point should have the greatest distance with the set $\widehat{\mathcal{G}}_{t}^{j}:=\{\widehat{\mathcal{G}}_{t}(\widetilde{m}_{i})\}_{i=1}^{j}$, i.e.,
$\widetilde{m}_{j+1}=\arg\max_{m\in\Gamma/\gamma^{j}}\left(d\big{(}\widehat{\mathcal{G}}_{t}(m),\widehat{\mathcal{G}}_{t}^{j}\big{)}-\lambda\|m-r^{t}\|_{2}\right),$ (11)